OpenStudy (anonymous):

Add or subtract as indicated. 7 - 5 A. 12 B. 2 C. -11 D. -3

OpenStudy (lilg132):

lol

OpenStudy (anonymous):

7 - 5 = 2 B

OpenStudy (anonymous):

7\[7\sqrt{7}-5\sqrt{7}\]

OpenStudy (lilg132):

take 5 away from 7 what do you get? (use your fingers to help)

OpenStudy (anonymous):

no it didnt work

OpenStudy (anonymous):

\(\sqrt7(7-5)\)

OpenStudy (anonymous):

2√7

OpenStudy (anonymous):

oops wrong question

OpenStudy (anonymous):

lol

OpenStudy (amistre64):



OpenStudy (anonymous):

\[\sqrt{3}(\sqrt{7} - \sqrt{5})\]

OpenStudy (anonymous):

mutiliply or simply

OpenStudy (anonymous):

do i distrubt the 3

OpenStudy (anonymous):

it's already simplified

OpenStudy (anonymous):

you can distribute but this form is ok too

OpenStudy (anonymous):

\[3\sqrt{7}+3\sqrt{5}\]

OpenStudy (anonymous):

it says simplify but this is right

OpenStudy (anonymous):

oh its suppose to be minus not add. So its \[\sqrt{27}-\sqrt{15}\]